OpenStudy (anonymous):

Write an equation of a line in slope-intercept form that contains the given points. 1. U(-2,-9) V(1,8)

OpenStudy (anonymous):

The Formula Is: y=mx+b

OpenStudy (katrinakaif):

First find the slope of both of them by using the formula y2-y1/x2-x1 Then use one of the orderp pairs and plug it in to the formula y-y1) = m(x-x1) using ur slope here which is m

OpenStudy (chaise):

y = mx+b y = (47/30)x+b (x,y) (1,8) 8 = (47/30)1+b b = 193/30 y = mx+b y = (47/30)x+(193/30) Hope this is right, did it really quickly so sorry if there is errors.

OpenStudy (chaise):

your m value was achieved by the equation: m = (y1-y2)/(x1-x2) if this helps.

OpenStudy (anonymous):

http://www.youtube.com/watch?v=u3spOO-m_Gg Here's A Link To Help :D